OpenStudy (anonymous):

Will give medal and fan to who ever helps me with this. Simplify: (y3)(y4)(y) = ?

OpenStudy (anonymous):

Please anyone! I really need help with this. I cant remember how to do it. :(

OpenStudy (anonymous):

@texaschic101

OpenStudy (texaschic101):

exponent law : (x^m)(x^n) = x^m+n) so... (y^3)(y^4) = y^7 and (y^7)(y) = y^8 ==> I believe this is your answer

OpenStudy (anonymous):

Ok that is what i was aiming for. But I just wanted to be sure. Thanks again!

OpenStudy (texaschic101):

anytime :)